OpenStudy (anonymous):

System of 3 variable Linear Equations a) 6x+18y-4z=-1, 11x+8y-9z=0, 3x-6y+0z=-7

OpenStudy (anonymous):

x = -482/225, y = 43/450, z = -38/15

OpenStudy (anonymous):

how do you do it?

OpenStudy (anonymous):

6 18 -4 -1 11 8 -9 0 3 -6 0 -7 6 18 -4 -1 3 -6 0 -7 11 8 -9 0 12 36 -8 -2 11 8 -9 0 3 -6 0 -7 1 28 1 -2 3 -6 0 -7 11 8 -9 0 1 28 1 -2 0 -90 -3 -1 0 -300 -20 22 you work out the rest... I give up

OpenStudy (anonymous):

How do you do the workings out? If you wrote it all out it would help a lot :)

OpenStudy (anonymous):

the working out is actually a procedure, or an algorithm :-D http://en.wikipedia.org/wiki/Gaussian_elimination better: http://en.wikipedia.org/wiki/Gauss%E2%80%93Jordan_elimination